STRUCTURE OF THE DENSE MOLECULAR GAS IN THE HELIX NEBULA: LARGE SCALE MAPPING OF HCO^{+}

\underline{N. R. ZEIGLER}, L. M. ZIURYS, Department of Chemistry, University of Arizona, PO Box 210041, 1306 East University Blvd, Tucson, AZ, 85721, USA; L. N. ZACK, Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.

The Helix Nebula, NGC 7293, has been mapped in HCO^{+} J = 1 → 0 emission with a 70′′ spatial resolution (1.68 km s^{−1} velocity resolution) using the Arizona Radio Observatory (ARO) 12 m telescope on Kitt Peak. Over 200 individual positions covering an area of roughly 1000 × 800 ′′ were observed down to a 3σ noise level of ˜ 20 mK. HCO^{+} was detected at over three-quarters of the positions, with the majority of spectra showing multiple velocity components indicative of a complex kinematic structure. The column density of HCO^{+} across the Helix ranges from N_{tot} ˜ 1.4 × 10^{11} to ˜ 2.4 × 10^{12} cm^{−2}, with an average N_{tot} ˜ 4 × 10^{11} cm^{−2}. The HCO^{+} distribution, which traces gas with densities ˜ 10^{5} cm^{−3}, is similar to that of CO and H_{2} , with certain point symmetries and red and blueshifted regions common in all three molecules; these data are also consistent with observations of atomic lines. The HCO^{+} emission appears to trace two ring-like structures emanating from the central star: one is blue-shifted and the other red-shifted with respect to the star, and possibly trace the edges of a bipolar outflow.